5-9(x-9)+12 whats the answer
Found 2 solutions by rfer, richard1234:
Answer by rfer(16322)

5-9x+81+12
-9x+98=0
-9x=-98
x=10.89

Answer by richard1234(7193)

Cannot be solved for x, since it's not an equation. We can simplify the expression to obtain

5 - (9x - 81) + 12

= 5 - 9x + 81 + 12

= -9x + 98